OpenStudy (gabylovesyou):

Identify the slope and y-intercept of the inequality 12x + 3y > 15. slope = 4, y-intercept = 5 slope = -4, y-intercept = 15 slope = -1/4, y-intercept = 5 slope = -4, y-intercept = 5

OpenStudy (anonymous):

12x + 3y > 15. 12x + 3y - 15 = 0 For y intercept = -c/b = - (-15)/3 = 5

OpenStudy (pradipgr817):

y>-4x+5 so d

OpenStudy (anonymous):

Slope = -a/b -12/3 = -4

OpenStudy (gabylovesyou):

so it is D! Thank you :)

OpenStudy (anonymous):

12x+3y>15 first consider 12x+3y=15 3y=15-12x y=15/3 - 12/3 *x y=5-4x m=-4 c=5 (y intercept)